OpenStudy (anonymous):

simplify the imaginary number √-8

OpenStudy (anonymous):

first you can divide out a 4. so now you have \[\sqrt{4\times-2 }\] Then you seperate the two and you have \[\sqrt{4}\times \sqrt{-2}\] do you know what to do next?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok, well next you simplify the four, so what is radical 4? 4 is a perfect square, remember.

OpenStudy (anonymous):

√4

OpenStudy (anonymous):

no, what is radical 4? in other words, what do you multiply by itself to get 4?

OpenStudy (anonymous):

can you plz just the answer just this time

OpenStudy (anonymous):

@DuckDynastyfan923

OpenStudy (anonymous):

just tell me plz

OpenStudy (anonymous):

ok, well radical 4 is 2. so your answer would be \[2\sqrt{-2}\]